OpenStudy (anonymous):

I need help setting this up... A function (f) is given by the formula...f(x)=A*e^(k*x) for constants (A) and (k) we also know that f(5)=2 and f(8)=5. Find numerical values for the constants A and k.

OpenStudy (swissgirl):

@dan815 Found it in closed questions :)

OpenStudy (swissgirl):

Would u use matrices

OpenStudy (dan815):

f(x)=A*e^(k*x) 2=A*e^(5k) k=(ln2/A )/5 plug and solve 5=A*e^(8k)

OpenStudy (swissgirl):

Okkk cool -.-

OpenStudy (swissgirl):

mordy is being annoying and banging on my keyboard lol

OpenStudy (swissgirl):

I should have known this ...

OpenStudy (anonymous):

what do i plug in?